# KnowGPT: Black-Box Knowledge Injection for Large Language Models

### 11 Dec 2023 v1

@@ 3.2.1 Implementation
  • $\mathcal{P}_{\text{sub}}$: A heuristic sub-graph extraction strategy that extracts a 2-hop subgraph around both the source and target entities. Detailed implementation can be found in Appendix A.1. Since RL is notoriously unstable Sutton & Barto (2018), we introduce $\mathcal{P}_{\text{sub}}$ as an alternative candidate strategy for the MAB selection, ensuring a fallback option if the RL-based approach does not perform well.
@@
Table 1: Performance comparison among baselines and KnowGPT on three benchmark datasets.
@@ 4.1.1 QA Datasets
We evaluate KnowGPT on three QA datasets spanning two fields: CommonsenseQA Talmor et al. (2019) and OpenBookQA Mihaylov et al. (2018) serve as benchmarks for commonsense reasoning, while MedQA-USMLE Jin et al. (2021) acts as a domain-specific QA benchmark. The statistics of these three datasets can be found in Table LABEL:tab:DatasetStatistics in the Appendix.
4.1.2 Background Knowledge
To facilitate common sense reasoning, we employ ConceptNet Speer et al. (2017), an extensive commonsense knowledge graph comprising more than 8 million interconnected entities through 34 concise relationships. For tasks specific to the medical domain, we leverage USMLE Yasunaga et al. (2021) as our foundational knowledge source. USMLE is a biomedical knowledge graph that amalgamates the Disease Database segment of the Unified Medical Language System (UMLS) Bodenreider (2004) and DrugBank Wishart et al. (2017). This repository encompasses 9,958 nodes and 44,561 edges.
@@ 4.1.3 State-of-the-art Baselines
We carefully select baseline models from three categories for a comprehensive evaluation.
LM + Fine-tuning We compare our method with vanilla fine-tuned LMs. Specifically, we choose Bert-base, Bert-large Devlin et al. (2018), and RoBert-large Liu et al. (2019) as representative fine-tune LM methods. To conduct commonsense and biomedical QA, we fine-tune these three LMs via an additional linear layer.
KG-enhanced LM We have also implemented several recently released models for integrating KGs into question answering, which encompass MHGRN Feng et al. (2020), QA-GNN Yasunaga et al. (2021), HamQA Dong et al. (2023a), JointLK Sun et al. (2021b), GreaseLM Zhang et al. (2022) and GrapeQA Taunk et al. (2023). To ensure a fair comparison, we implement these baselines with advanced language models that are optimized for particular datasets. Specifically, RoBert-large Liu et al. (2019) is used for CommenseQA, while AristoRoBERTa Clark et al. (2020) is designated for OpenBookQA. For MedQA, we opt for the top-tier biomedical language model, SapBERT Liu et al. (2020a). Note that due to the white-box nature of these methods and their high computation overheads, it is infeasible to apply them to state-of-the-art LLMs, like ChatGPT and GPT-4.
LLM We also add several representative generative LLMs, including ChatGLM, ChatGLM2, Baichuan-7B, InternLM, GPT-3, ChatGPT and GPT-4 as knowledge-agnostic alternatives. Specifically, we used the model ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of ChatGPT and GPT-4, respectively (we have provided more implementation details of all LLMs in Appendix A.4). The question-answering task is conducted under the zero-shot setting with the question query from the test set as input.
@@ 4.2 Main Results (RQ1)
To address RQ1, we evaluate KnowGPT by comparing it to state-of-the-art baselines on the three benchmark datasets. KnowGPT is based on the original ChatGPT. We measure the performance using accuracy, which calculates the percentage of questions correctly predicted by the model out of the total questions in the test set. We make the following observations:
@@
  • KnowGPT outperforms all KG-enhanced LMs significantly. This implies our black-box knowledge injection method proficiently encodes knowledge into LLMs. Furthermore, it showcases the superiority of our black-box approach, given its adaptable application to ChatGPT using only the model API, a feat not achievable by white-box methods.
@@ 4.2.1 Leaderboard Ranking
We submit our testing results on the official leaderboard maintained by the authors of OpenbookQA. The full records on the leaderboard are shown on the website [^1] https://leaderboard.allenai.org/open_book_qa/submissions/public. , while our result can be found from here [^2] https://leaderboard.allenai.org/open_book_qa/submission/cj9game4arcuacugbrj0..
We summarize the related submissions in Table 2, including three categories: traditional KG-enhanced LM, fine-tuning of LLMs, e.g., T5-11B used in UnifiedQA, and ensemble of multiple predictors. KnowGPT significantly outperforms traditional KG-enhanced LMs with 4.2% improvements when compared to the best baseline. The third group of methods occupies the leaderboard by leveraging ensemble learning. Nevertheless, KnowGPT can still obtain competitive performance without ensembling with merely 0.4% below GenMC Ensemble Huang et al. (2022). Additionally, our KnowGPT is also comparable to human performance.
@@
4.3 Ablation Studies (RQ2)
To answer RQ2, we conduct two ablation studies. First, in Table 3, we measure the importance of the tailored reinforcement learning-based path extraction module, i.e., $\mathcal{P}_{\text{RL}}$. Specifically, we compare it with the direct path extraction method, i.e., $\mathcal{P}_{\text{sub}}$. The performance is evaluated by directly feeding the extracted knowledge with the prompt format of ‘Sentence’, i.e., $\mathcal{F}_{s}$, to ChatGPT. We also include ‘w/o KG’ as the baseline where ChatGPT is asked to independently answer the given question with no reasoning background provided. The results clearly indicate the vital role of our proposed path extraction strategies. Second, we compare each of the three prompt formats subject to the same extracted knowledge. The detailed results are shown in Table 4. Though different formats perform similarly within the difference of 2.2% - 3.3%, they are particularly suitable for different kinds of questions. We illustrate this observation in the following case study section. Both ablation studies support the indispensability of each module, armed with a tailored deep reinforcement learning-based path extraction and a context-aware prompt translation, our KnowGPT performs best on all three benchmark datasets.
Table 4: Ablation study on the effectiveness of prompt translation formats.
@@ 4.3 Ablation Studies (RQ2)
4.4 Case Studies (RQ3)
@@
5 Related Work
5.1 Integration of KGs and LLMs
One heuristic way to incorporate KGs and LLMs is to inject triples as the input Pan et al. (2023). ERNIE3.0 Sun et al. (2021a) takes triples as token sequences and straightforwardly appends them to the given sentences. Colake Sun et al. (2020) presents a unified graph that combines a word graph with the given context and KG. QA-GNN Yasunaga et al. (2021) takes this further and embeds the question context as an entity in the joint graph. Recent studies utilize the attention mechanism to incorporate KGs into LMs, thereby enhancing comprehension and reasoning processes. HamQA Dong et al. (2023a) proposes a hyperbolic-based graph attention network to learn from the ubiquitous hyponymy in real-world questions. However, these methods assume they know everything about LMs, including model architectures and parameters. However, many SOTA LLMs are confined to a black-box role in practice. In this work, we develop a black-box knowledge injection framework that can efficiently integrate KGs into LLMs with APIs only.
5.2 Prompt-based Learning
Prompting texts to LLMs for prediction has been a new learning paradigm in natural language processing Liu et al. (2023). In this paper, we refer to fixed-LM prompt tuning Li & Liang (2021); Lester et al. (2021), where the model only optimizes the prompt construction and keeps the LM frozen. To form effective prompts with suitable templates, researchers propose to shape the contexts into clozes with unfilled slots Jiang et al. (2020); Petroni et al. (2019), or manually define the prompt format for different scenarios Brown et al. (2020). In this paper, we propose a knowledge injection framework that $(i)$ replaces the cloze and examples with knowledgeable reasoning background and $(ii)$ employs a tailored Multi-armed Bandit algorithm to automatically select the suitable prompt formats.
6 Conclusion and Future Work
In this work, we formally define the problem of black-box knowledge injection for LLMs in complex question answering. A novel framework, namely KnowGPT, is presented to integrate KGs into LLMs effectively with model APIs only. We first train a deep RL policy to extract informative and concise reasoning paths from the KG. Then we learn an MAB to select the most effective path extraction method and prompt template for each question. Extensive experiments on both general and domain-specific QA datasets show the superior performance of KnowGPT and the effectiveness of each component. In the future, we will study more advanced path extraction strategies and prompt templates to improve the performance further.

### 13 Mar 2024 v4

@@ 4.2.1 Implementation
  • $\mathcal{P}_{\text{sub}}$: A heuristic sub-graph extraction strategy that extracts a 2-hop subgraph around both the source and target entities. Detailed implementation can be found in Appendix A.1. Since RL is notoriously unstable Sutton and Barto (2018), we introduce $\mathcal{P}_{\text{sub}}$ as an alternative candidate strategy for the MAB selection, ensuring a fallback option if the RL-based approach does not perform well.
@@
Table 1: Performance comparison among state-of-the-art baselines and KnowGPT on three benchmark datasets.
@@
*The results compared with fine-tuning LLMs on CommonsenseQA are placed in Table 7 of Appendix.
@@ 5.1.1 QA Datasets
We evaluate KnowGPT on three QA datasets spanning two fields: CommonsenseQA Talmor et al. (2019) and OpenBookQA Mihaylov et al. (2018) serve as benchmarks for commonsense reasoning, while MedQA-USMLE Jin et al. (2021) acts as a domain-specific QA benchmark. The statistics of these three datasets can be found in Table 5 in the Appendix.
@@ 5.1.2 State-of-the-art Baselines
We carefully select baseline models from four categories for a comprehensive evaluation.
LM + Fine-tuning We compare our method with vanilla fine-tuned LMs. Specifically, we choose Bert-base, Bert-large Kenton and Toutanova (2019), and RoBerta-large Liu et al. (2019) as representative fine-tune LM methods. To conduct commonsense and biomedical QA, we fine-tune these three LMs via an additional linear layer.
KG-enhanced LM We have also implemented several recently released models for integrating KGs into question answering, which encompass MHGRN Feng et al. (2020), QA-GNN Yasunaga et al. (2021), HamQA Dong et al. (2023a), JointLK Sun et al. (2022), GreaseLM Zhang et al. (2022b) and GrapeQA Taunk et al. (2023). To ensure a fair comparison, we implement these baselines with advanced language models that are optimized for particular datasets. Specifically, RoBerta-large Liu et al. (2019) is used for CommenseQA, while AristoRoBERTa Clark et al. (2020) is designated for OpenBookQA. For MedQA, we opt for the top-tier biomedical language model, SapBERT Liu et al. (2021). Note that due to the white-box nature of these methods and their high computation overheads, it is infeasible to apply them to state-of-the-art LLMs, like ChatGPT and GPT-4.
LLM We include several representative generative LLMs, including ChatGLM, ChatGLM2, Baichuan-7B, InternLM, GPT-3, ChatGPT and GPT-4 as knowledge-agnostic alternatives. Specifically, we used the model ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of ChatGPT and GPT-4, respectively (we have provided more implementation details of all LLMs in Appendix A.4). The question-answering task is conducted under the zero-shot setting with the question query from the test set as input.
LLM + KG Prompting To verify the effectiveness of our proposed prompting strategy used KnowGPT, we also add the state-of-the-art KG prompting methods, i.e., CoK Wang et al. (2023), ChatRule Luo et al. (2023), and Mindmap Wen et al. (2023) as baselines.
@@ 5.2 Main Results (RQ1)
To address RQ1, we evaluate KnowGPT by comparing it to state-of-the-art baselines on the three benchmark datasets. KnowGPT is based on the original ChatGPT. We measure the performance using accuracy, which calculates the percentage of questions correctly predicted by the model out of the total questions. We have following observations:
@@
  • KnowGPT outperforms all KG-enhanced LMs significantly. This implies our black-box knowledge injection method proficiently encodes knowledge into LLMs. Furthermore, it showcases the superiority of our model, given its adaptable application to ChatGPT using only the model API, a feat not achievable by white-box methods.
@@ 5.2.1 Leaderboard Ranking
We submit our results onto the official leaderboard maintained by the authors of OpenbookQA. The full records on the leaderboard are shown on the website [^1] https://leaderboard.allenai.org/open˙book˙qa/submissions/public. , while our result can be found from here [^2] https://leaderboard.allenai.org/open˙book˙qa/submission/cj9game4arcuacugbrj0..
We summarize the related submissions in Table 2, including three categories: traditional KG-enhanced LM, fine-tuning of LLMs, e.g., T5-11B used in UnifiedQA, and ensemble of multiple predictors. KnowGPT significantly outperforms traditional KG-enhanced LMs with 4.2% improvements when compared to the best baseline. Despite the 2.6% difference of the state-of-the-art fine-tuning method X-Reasoner, which is the currently best single model, they costly train the T5-11B LLM and also initialize the model on eight A100 GPUs. The third group of methods occupies the leaderboard by leveraging ensemble learning strategies. Nevertheless, KnowGPT can still obtain competitive performance without ensembling with merely 0.4% below GenMC Ensemble Huang et al. (2022). Notably, our KnowGPT is remarkably comparable to the human performance.
@@
Table 4: Ablation study on different prompt translation formats for the extracted knowledge.
@@
5.3 Ablation Studies (RQ2)
To answer RQ2, we conduct two ablation studies. First, in Table 3, we measure the importance of the tailored reinforcement learning-based path extraction module, i.e., $\mathcal{P}_{\text{RL}}$. Specifically, we compare it with the heuristic sub-graph extraction strategy, i.e., $\mathcal{P}_{\text{sub}}$. The performance is evaluated by directly feeding the extracted knowledge with the prompt format of ‘Sentence’, i.e., $\mathcal{F}_{s}$, to ChatGPT. We also include ‘w/o KG’ as the baseline where ChatGPT is asked to independently answer the given question with no reasoning background provided. The results clearly indicate the vital role of our proposed path extraction strategies. Second, we compare each of the three prompt formats subject to the same extracted knowledge. The detailed results are shown in Table 4. Though different formats perform similarly within the difference of 2.2% - 3.3%, they are particularly suitable for different kinds of questions. We illustrate this observation in the following case study section. Both ablation studies support the indispensability of each module, armed with a tailored deep reinforcement learning-based path extraction and a context-aware prompt translation, our KnowGPT performs best on all three benchmark datasets.
6 Conclusion and Future Work
In this work, we formally define the problem of knowledge injection for LLMs in complex question answering. A novel framework, namely KnowGPT, is presented to integrate KGs into LLMs effectively with model APIs only. We first train a deep RL policy to extract informative and concise reasoning paths from the KG. Then we learn an MAB to select the most suitable and effective path extraction method and prompt template subject to each question. Extensive experiments on both general and domain-specific QA datasets show the superior performance of KnowGPT and the effectiveness of each component. In the future, we will study more advanced path extraction strategies and prompt templates to improve the performance further.
Limitations
Through our exploration, we realize the natural limitations of KnowGPT brought by real-world KGs. Existing KGs are automatically constructed based on online corpora. This inevitably introduces a considerable number of noisy triples into KGs. Under such circumstances, the noisy knowledge may mislead the LLMs to wrong predictions despite the effectiveness of our injection methods. To apply KnowGPT to practical scenarios, we would leverage off-the-shelf KG refinement algorithms to improve the quality of KGs.
Ethics Statement
In this study, all the datasets are publicly available and have been extensively used in research related to neural language processing.
Appendix A Case Studies (RQ3)
@@
Appendix B Implementation Details
B.1 Entity Linking and Heuristic Path Extraction $\mathcal{P}_{sub}$
For each QA context, we adopt the methodology outlined in the prior research Lin et al. (2019); Yasunaga et al. (2022) to extract the subgraph from the background knowledge graph (KG), denoted as $\mathcal{G}$. We commence by executing entity linking on $\mathcal{G}$, resulting in an initial collection of nodes, $V_{topic}$. Next, we incorporate bridge entities that appear within a 2-hop path between any two linked entities from $V_{topic}$, yielding the set $V_{retrieval}$. Subsequently, we refine this set by evaluating the relevance score for each node, adhering to the method proposed Yasunaga et al. (2022). From this refined set, only the top 200 nodes, based on score, are retained, and the others are discarded. We then extract all edges connecting any pair of nodes in $V_{sub}$, creating the retrieved subgraph $G_{sub}$. Each node within $G_{sub}$ is designated a type based on its association to either the topic entities $Q$ or target entities $A$. Intuitively, the relevant reasoning background lies in a question-specific subgraph $\mathcal{G}_{sub}$ that contains all the source entities $S$, target entities $A$, and their $k$-hop neighbors. Therefore, the reasoning background could be provided as the $\mathcal{G}_{sub}$, we denote this direct path extraction method as $\mathcal{P}_{sub}$.
B.2 Graph Initialization
To calculate the initial node embeddings for entities extracted from the background KG, we adopt the approach proposed by the previous study Feng et al. (2020). Specifically, we transform knowledge triples from the KG into sentences and feed them into pre-trained LMs to get node embeddings. Specifically, to ensure a fair comparison, we implement all the KG-enhanced baselines and our model with the same advanced language models that are optimized for particular datasets. Specifically, RoBert-large Liu et al. (2019) is used for CommenseQA, while AristoRoBERTa Clark et al. (2020) is designated for OpenBookQA. For MedQA, we opt for the top-tier biomedical language model, SapBERT Liu et al. (2021), to enhance comprehension of the biomedical field.
B.3 Datasets
We evaluate KnowGPT on three QA datasets spanning two fields: CommonsenseQA Talmor et al. (2019) and OpenBookQA Mihaylov et al. (2018) serve as benchmarks for commonsense reasoning, while MedQA-USMLE Jin et al. (2021) acts as a domain-specific QA benchmark. While the official test set serves primarily for leaderboard rankings, we initially assess model efficacy using the in-house (IH) data split introduced in Lin et al. (2019). The official dataset is denoted as CSQA, while the IH split is represented by CSQA(IH)*. The statistics of these three datasets can be found in Table 5.
Table 5: The statistical information of three datasets.
Dataset Question Choices Train Dev Test
CSQA #12102 5 $9741$ $1221$ $1140$
CSQA(IH) #12102 5 $8500$ $1221$ $1241$
OBQA #5957 4 4957 500 500
MedQA #12723 4 10178 1272 1273
@@ B.3 Datasets
Background Knowledge To facilitate common sense reasoning, we employ ConceptNet Speer et al. (2017), an extensive commonsense knowledge graph comprising more than 8 million interconnected entities through 34 concise relationships. For tasks specific to the medical domain, we leverage USMLE Yasunaga et al. (2021) as our foundational knowledge source. USMLE is a biomedical knowledge graph that amalgamates the Disease Database segment of the Unified Medical Language System (UMLS) Bodenreider (2004) and DrugBank Wishart et al. (2017). This repository encompasses 9,958 nodes and 44,561 edges.
B.4 Implementation of Baselines
To verify the effectiveness of our proposed KnowGPT, we carefully selected baseline models from three aspects to ensure a comprehensive evaluation, among which Bert-base, Bert-large Kenton and Toutanova (2019), and RoBert-large Liu et al. (2019) are picked for being representative fine-tune LM methods; MHGRN Feng et al. (2020), QA-GNN Yasunaga et al. (2021), HamQA Dong et al. (2023a), JointLK Sun et al. (2022), GreaseLM Zhang et al. (2022b) and GrapeQA Taunk et al. (2023) represent the state-of-art KG-enhanced LMs; ChatGLM Du et al. (2022), ChatGLM2 Zeng et al. (2022), Baichuan-7B, InternLM Team (2023), GPT-3 Brown et al. (2020), ChatGPT Ouyang et al. (2022) and GPT-4 OpenAI (2023) are picked for being representative generative large language models. Notably, while some LLM baselines are actually open-source, we conduct the question-answering task under the zero-shot setting with the question query from the test set as input. All baseline methods used in this paper are based on their open-source implementations or officially-released APIs. We list the source links in Table 6. Notably, we used the model ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of ChatGPT and GPT-4, respectively.
Table 6: Implementation codes for baselines.
Appendix C Subplementary Results
C.1 Fine-tune LLMs
To further verify the effectiveness of the knowledge injection framework, we also add several open-source trainable LLMs, i.e., ChatGLM, ChatGLM2, LLaMA-B, Baichuan-7B, InternLM and Vicuna-7B, and fine-tune them for commonsense reasoning on the benchmark CommonseQA. As shown in the Table 7, Our KnowGPT achieves comparable performance with no tuning on the LLM.
Table 7: Fine-tuned LLMs on CommonsenseQA.
LLM Acc
ChatGLM 0.559
ChatGLM2 0.600
LLaMA-7B 0.650
Baichuan-7B 0.588
Alpaca-7B 0.687
Vicuna-7B 0.667
InternLM-7B 0.752
KnowGPT 0.818
C.2 The effect of prompt format on different types of questions.
In this part, we conduct comprehensive experiments to investigate the effect of prompt format on different types of questions. Table 8 presents the accuracy of different prompt formats on three types of questions. We observe that graph description-based prompt performs significantly better than any other prompt formats on complex questions. It is because graph description-based prompt could provide LLMs with more detailed and structured information by highlighting the local structure of the central entity.
Table 8: Accuracy of different prompt formats on specific types of questions on CommonsenseQA.
Prompt Format Simple Multi-hop Graph reasoning
$\mathcal{F}_{t}$ (Triple) 0.9412 0.7447 0.4210
$\mathcal{F}_{s}$ (Sentence) 0.8823 0.8298 0.4739
$\mathcal{F}_{g}$ (Graph) 0.8529 0.7021 0.7895